1/3 divided by2/5 =6/x can you help me solve it?
Answer by ptaylor(1894)

1/3 divided by2/5 =6/x can you help me solve it?
Sure.
(1/3)/(2/5)=6/x
First Let's work on the left side. We'll multiply the left side by 5/5. That should simplify the fraction on the left.
5(1/3)/5(2/5)=(5/2)(1/3)=5/6 Now we have:
5/6=6/x Now we will multiply both sides by 6x:
6x(5/6)=6x(6/x) and we have:
5x=36
x=7.2
ck
.33333/.4=6/7.2
.83333=.833333
Hope this helps----ptaylor